# Hénon maps with many rational periodic points

### v1

@@ 3.2. Tail behaviour of $s_{d}$
$+\frac{1}{(d+2)!}x(x-1)\cdots(x-d-1)\left(-\frac{2x-(d+1)}{(d+3)(d+4)}\right).$
@@
  • By considering its value at $x=d+3$,
    $1-\frac{(x+1)(x-d-2)}{(d+3)(d+4)}\geq 1-\frac{(d+4)\cdot 2}{(d+3)(d+4)}=\frac{d+1}{d+3}.$
  • By considering $x=d+3$,
    $\sum_{j=0}^{d+1}\prod_{i\neq j,0\leq i\leq d+1}(x-i)$ $\geq(d+3)!\left(\frac{1}{2}+\frac{1}{3}+\cdots+\frac{1}{d+3}\right)$
    $=(d+3)!(H_{d+3}-1)$
    $>(d+3)!(\log(d+3)-1)$
    $>(d+3)!\log d.$
  • By considering $x=d+4$,
    $\frac{1}{(d+2)!}x(x-1)\cdots(x-d-1)\leq\frac{1}{(d+2)!}(d+4)(d+3)\cdots 3=\frac{(d+4)(d+3)}{2}.$
  • By considering $x=d+4$,
    $\frac{2x-(d+1)}{(d+3)(d+4)}\leq\frac{2(d+8)-(d+1)}{(d+3)(d+4)}=\frac{d+7}{(d+3)(d+4)}<\frac{2}{d+3}.$
@@
$\tilde{c}_{d+2}^{\prime}(x)-\tilde{c}_{d+4}^{\prime}(x)$ $>\frac{1}{(d+2)!}\frac{d+1}{d+3}(d+3)!\log d-\frac{(d+4)(d+3)}{2}\frac{2}{d+3}$
$>(d+1)\log d-d-4>\frac{2\pi}{3\sqrt{3}}+1.$
on $d+3\leq x\leq d+4$ for all $d\geq 10$. This proves the lemma. ∎
@@ 4. Hénon maps with many periodic points
for odd $d\geq 3$. This family exhibits many interesting properties, namely having a large number of periodic integer points near the origin with long periodic cycles. We focus on the case of $d$ odd to minimize tedium and simplify estimates: it is likely that similar results hold for some suitably chosen shift in the case of $d$ even.
@@
where $r(x,y)=(y,x)$ is a map that reflects the $x$ and $y$ coordinates. Since $d$ is odd, $h_{d}$ and its inverse preserve the lattice $\mathbb{Z}^{2}$.
@@ 4.1. Escape radius
For each $v\in S$, consider the sets:
@@ 5.2. Transcendental Hénon maps
Recall from Proposition 3.6 that $(-1)^{d}s_{2d+1}(y)$ locally uniformly converges to $\frac{2}{\sqrt{3}}\sin(\frac{\pi x}{3})$ on the real line as $d$ tends to infinity. We conclude this article by studying the dynamical behaviour of the limiting Hénon map:

### v2

@@ 3.2. Tail behaviour of $s_{d}$
$-\frac{1}{(d+2)!}x(x-1)\cdots(x-d-1)\left(\frac{2x-d-1}{(d+3)(d+4)}\right).$
@@
  • By putting $x=d+4$,
    $1-\frac{(x+1)(x-d-2)}{(d+3)(d+4)}\geq 1-\frac{2(d+5)}{(d+3)(d+4)}\geq\frac{d}{d+3}.$
  • By putting $x=d+3$,
    $\sum_{j=0}^{d+1}\prod_{i\neq j,0\leq i\leq d+1}(x-i)$ $\geq(d+3)!\left(\frac{1}{2}+\frac{1}{3}+\cdots+\frac{1}{d+3}\right)>(d+3)!\left(\log d-1/2\right),$
    where we use the lower bound $H_{d}>\log d+1/2$.
  • By putting $x=d+4$,
    $\frac{1}{(d+2)!}x(x-1)\cdots(x-d-1)\leq\frac{(d+4)(d+3)}{2}.$
  • By putting $x=d+4$,
    $\frac{2x-d-1}{(d+3)(d+4)}\leq\frac{d+7}{(d+3)(d+4)}<\frac{2}{d+3}.$
@@
$\tilde{c}_{d+2}^{\prime}(x)-\tilde{c}_{d+4}^{\prime}(x+1)$ $>d\left(\log d-1/2\right)-d-4$
on $d+3\leq x\leq d+4$. Since $d\left(\log d-1/2\right)-d-4>\frac{2\pi}{3\sqrt{3}}+2$ for all $d>44$, this proves the lemma. ∎
@@
We conclude this subsection with a lower bound of $s_{d}$ on some interval in the integers. This result will be useful in proving the boundedness of $K(r_{d})$ and understanding the preperiodic points of $r_{d}$, and will also be used in Section 4 when analyzing the periodic points of the Hénon map $h_{d}$.
Lemma 3.6.
Proof.
We proceed by induction on $d$. The claim is true for $s_{3}(x)=\frac{x^{3}-7x}{6}$. Now take $d\geq 4$ and suppose claim is true for some $d-1$. Take $x=\frac{d+7+2k}{2}$ with integer $k\geq 0$, then by $s_{d-1}=\delta s_{d}$ and the induction hypothesis we obtain
$s_{d}(x)$ $=s_{d}\left(\frac{d+5}{2}\right)+\sum_{j=0}^{k}s_{d-1}\left(\frac{d+6+2j}{2}\right)$
$\geq s_{d}\left(\frac{d+5}{2}\right)+\sum_{j=0}^{k}\frac{3}{2}(d+6+2j)=s_{d}\left(\frac{d+5}{2}\right)+\frac{3}{2}(k+1)(d+6+k).$
Since $s_{d}(\frac{d+5}{2})\geq d+1$ by Lemma 4.2(5) of [doyle2022polynomialsrationalpreperiodicpoints], we have
$s_{d}(x)-3x$ $\geq d+1+\frac{3}{2}(k+1)(d+6+k)-\frac{3}{2}(d+7+2k)$
$=d-\frac{1}{2}+\frac{3}{2}k(k+d+5)\geq\frac{9}{2}+\frac{3}{2}k(k+9)>0$
for all $k\geq 0$ and $d\geq 4$. Therefore the claim is true for $d$. ∎
@@ 4. Hénon maps with many periodic points
for odd $d\geq 3$. Recall that $s_{d}$ is integer-valued for odd $d$, so $h_{d}$ (and its inverse) preserves the integral lattice $\mathbb{Z}^{2}$. This family exhibits many interesting properties, namely having a large number of periodic integer points near the origin with long periodic cycles.
We are interested in asymptotic results so focus on the case of $d$ odd to simplify the estimates and exposition: similar results will likely hold for some suitably chosen shift in the case of $d$ even.
@@
where $r(x,y)=(y,x)$ is a map that reflects the $x$ and $y$ coordinates.
@@ 4.1. Escape radius
For each rational place $v$, consider the sets:
@@ 5.2. Transcendental Hénon maps
Recall from Proposition 3.1 that $(-1)^{d}s_{2d+1}(x)$ locally uniformly converges to $\frac{2}{\sqrt{3}}\sin(\frac{\pi x}{3})$ on the real line as $d$ tends to infinity. We conclude this article by studying the dynamical behaviour of the limiting Hénon map: